pooja195 (pooja195):

An office has two envelope stuffing machines. Machine A can stuff a batch of envelopes in 5 hours, while Machine B can stuff a batch of envelopes in 3 hours. How long would it take the two machines working together to stuff a batch of envelopes?

OpenStudy (anonymous):

hello @pooja195

OpenStudy (anonymous):

hmm... let's see what I can do: given: machine A: 1 batch/ 5 hours machine B: 1batch/ 3 hours together to make 1 batch: 1 batch= (rate of A)(time spent) + (rate of b)(time spent) 1= 1/5t + 1/3t evaluate and solve for t (time spent together)

pooja195 (pooja195):

i would do the reciprocal right?

OpenStudy (anonymous):

i'm not sure if it is necessary though first simplify the two terms on the right side, then let's see what's next

pooja195 (pooja195):

im think 3t+5t=15 8t=15 t=15/8?

OpenStudy (anonymous):

oh wow, lol i guess your right ^_^ got the same oh yeahhh, now I get it... reciprocal of 1 is 1, so yes we can do that

pooja195 (pooja195):

LOL xD ok So, it looks like it will take the two machines, working together, 1.875 hours to stuff a batch of envelopes.

OpenStudy (anonymous):

eggsactly :D

pooja195 (pooja195):

yay!!! thanks :)

OpenStudy (anonymous):

np! :P